Book A weighs 1.2 pounds and Book B weighs 1.3 pounds. If there are a total of 20 of these books in a box and the total weight of the books is 25.3 pounds, how many of each type of book are in the box? I set up A+B=20; B=20-A; and 1.2A+1.3B-25.3. But I can not figure out from there, please advise.
Answer by ikleyn(53765)

.
Book A weighs 1.2 pounds and Book B weighs 1.3 pounds. If there are a total of 20 of these books in a box and
the total weight of the books is 25.3 pounds, how many of each type of book are in the box?
I set up A+B=20; B=20-A; and 1.2A+1.3B-25.3.
But I can not figure out from there, please advise.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

   A +    B = 20,     (1)
1.2A + 1.3B = 25.3.   (2)

From (1), B = 20-A. Substitute it into (2). You will get

1.2A + 1.3*(20-A) = 25.3,

1.2A + 26 - 1.3A = 25.3,

-0.1A = 25.3 - 26,

-0.1A = -0.7.


Can you continue and complete?